OpenStudy (anonymous):

18. A customer went to a garden shop and bought some potting soil for $12.50 and 6 shrubs. If the total bill was $71.00, how much did each shrub cost if they all cost the same? a) Choose and label a variable for this situation. b) Write an equation using the variable that would answer the question. c) Solve your equation, and check your answer. d) Use the solution to write a sentence answer to the question in the word problem.

OpenStudy (lilg132):

71 - 12.50 = 58.50 58.50 is the cost of 6 shrubs to work out the price of each shrub divide this by 6 58.5 / 6 = 9.75 remember to click good answer

OpenStudy (anonymous):

did you sorl all the work?

OpenStudy (lilg132):

6x = 71 - 12.50 6x = 58.50 6x / 6 = 58.50 / 6 x = 9.75 each shurb costs 9.75

OpenStudy (lilg132):

x = shrubs

OpenStudy (hoblos):

a) x b) 6x +12.5 = 71 c) x=9.75 6*9.75 + 12.5 = 71 d)each shrub costs $9.75